A solenoid has 2000 turns on its 40cm length. What is the magnetic field intensity at its centre if it carries a current of 5A


1
Expert's answer
2022-01-24T08:45:08-0500

The magnetic field at the centre of such a solenoid can be found as


"B=\mu \dfrac NLI,\\\space\\
B=4\pi·10^{-7}·\dfrac{2000}{0.4}·5=0.0314\text{ T}."